OpenStudy (pandasneakers):

Help PLz The table shows the outputs y for different inputs x: Input (x) 5 6 7 8 Output (y) 2 5 8 11 Part A: Do the data in this table represent a function? Justify your answer. (3 points) Part B: Compare the data in the table with the relation f(x) = 2x + 13. Which relation has a greater value when x = 7? (2 points) Part C: Using the relation in Part B, what is the value of x if f(x) = 75? (5 points)

OpenStudy (dayakar):

A) it represents a function all the first co- ordinates( x values) are different

OpenStudy (pandasneakers):

@dayakar can u help with b this is what i have > fx=2x+13 7=2(7)+13 7=17 is this correct or not ?

OpenStudy (pandasneakers):

or do i imput the other numbers till i get 7 ?

OpenStudy (pandasneakers):

@ShelbyWyatt can u help plz?

OpenStudy (shelbywyatt):

lets see

OpenStudy (welshfella):

y = 2x + 13 gives 27 when x = 7 and the first function gives 8 when x = 7

OpenStudy (shelbywyatt):

i dont think i learned this yet

OpenStudy (welshfella):

- so 2x + 13 gives the greater value.

OpenStudy (welshfella):

for Part C you need to solve the equation 75 = 2x + 13 ( f(x) is given as = 75)

OpenStudy (shelbywyatt):

i get it now

OpenStudy (pandasneakers):

x=31 ?

OpenStudy (skullpatrol):

Correct :-)

OpenStudy (pandasneakers):

and thank you for all your help :D

OpenStudy (skullpatrol):

Thanks for trying to learn.

OpenStudy (welshfella):

yes its refreshing to see someone trying to learn rather than just wanting the answer.